Grade Upto college levelMechanics

a motor drives a body along a straight line with a constant force .the power P developed must vary with time as ......

plzzz plot a power vs time graph for this with explanation ......................................... i will surely approve.

iit jee
15 Years agoGrade Upto college level

1 Answer

vikas askiitian expert
ApprovedApproved Tutor Answer15 Years ago

force = F = constant

 

ma = F

a = F/m

 

v  = u + at  

u = 0 so

v = at = Ft/m          .................1

 

power = force*velocity = F2t/m

power = kt

power is directly proportional to time , graph bw power & t is straight line .....